wilson has 8 coins of same value and seven paise. Chloe has 11 coins of same value as those that wilson has and she has five paise. Wilson says "If you give me one of your coins and four paise, will have same amount of your money". What is the value of each coin that wilson and chloe have?
Answer by greenestamps(13200)



Let x be the value of the unknown coins. Then

8x+7 = paise Wilson has
11x+5 = paise Chloe has

If Chloe gives Wilson one coin and 4 paise, then

(8x+7) + (x+4) = 9x+11 will be the paise Wilson has
(11x+5) - (x+4) = 10x+1 will be the paise Chloe has

Those two amounts will be the same

9x+11 = 10x+1
x=10

The value of the coins that Wilson and Chloe have is

ANSWER: 10 paise